# Can Large Language Models Improve Phishing Defense? A Large-Scale Controlled Experiment on Warning Dialogue Explanations

### v1

@@ Appendix A Warning Messages
In the following are reported the explanation messages contained in the warnings for each condition. Each condition had three warnings, one per email: IP address, Top-level domain mispositioned, and Link mismatch.
A.1 Claude feature-based explanations
  • 1.
@@ A.1 Claude feature-based explanations
A.2 Llama feature-based explanations
  • 1.
@@ A.2 Llama feature-based explanations
  • 3.
    Link mismatch. The link “https://www.facebook.com/hacked/disavow? u=100000125023309&nArdInDS2&lit_IT&ext1548538159” is an imitation of the original Facebook link. This site might be intended to take you to a different place. You might be disclosing private information.
A.3 Claude counterfactual explanations
  • 1.
@@ A.3 Claude counterfactual explanations
A.4 Llama counterfactual explanations
  • 1.
@@ A.4 Llama counterfactual explanations
  • 3.
    Link mismatch. The link text “protect your account” hides the actual link “https://www.facebook.com/hacked/disavow?u=100000125023309 &nArdInDS2&lit_IT&ext1548538159” The email would have been considered safe if the link text matched the actual link, which helps ensure the link is trustworthy and not trying to trick you. A safe link might have looked like: https://www.facebook.com/account-security.
A.5 Baseline - Manually-generated feature-based explanations
  • 1.
@@ A.5 Baseline - Manually-generated feature-based explanations
  • 2.
    Top-level domain mispositioned. In the URL present in the email (“https://amazonservices.com.cz/ account.php”), the top-level domain (e.g., “.com“) is in an abnormal position. This could indicate that the URL leads to a fake website. Such websites might steal your personal information.
  • 3.
    Link mismatch. This email reports a link that is different from the actual one “https://www.facebook.com/hacked/disavow? u=1000001250 23309&nArdInDS2&lit_IT&ext1548538159”. This site might be intended to take you to a different place. You might be disclosing private information.
A.6 Logistic Regression Detailed Results
Table 4: Logistic regression results on CTR. OR¿1 increases click likelihood; OR¡1 reduces it. Effect sizes: Small (OR 1.2–1.5/0.67–0.83), Moderate (1.51–2/0.5–0.66), Large (¿2/¡0.5). ORs between 0.84 and 1.19 are reported as Negligible.
Variable Condition Emails OR $p$ Effect
@@ A.6 Logistic Regression Detailed Results
Action=“don’t continue” CF agg ALL 0.21 ¡.001 Large
CF agg TP 0.18 ¡.001 Large
NoLLM ALL 0.14 ¡.001 Large
NoLLM TP 0.10 ¡.001 Large
@@
Avg hours online/day NoLLM ALL 1.24 .024 Small
@@
Age: Avg hours online/day FB agg ALL 1.007 .025 Negligible
CF agg ALL 1.007 .025 Negligible
Llama agg ALL 1.007 .025 Negligible
Claude agg ALL 1.007 .025 Negligible
NoLLM TP 1.02 .027 Negligible
Claude CF FP 0.93 .014 Negligible
Expertise:Avg hours online/day Claude FB ALL 1.07 .042 Negligible
Llama CF ALL 1.11 .039 Negligible
Gender=Male: Need-for-Cognition CF agg FP 1.44 .001 Moderate
@@
Llama agg ALL 1.04 .032 Negligible
Expertise:Need-for-Cognition Llama agg ALL 1.03 .032 Negligible
Age:Expertise CF agg ALL 0.98 .048 Negligible
CF agg FP 0.97 .045 Negligible
Gender=Male:Age Llama agg TP 0.88 .02 Negligible

### v2

@@ 9.2 Explanation generation prompts
Below are the explanation messages contained in the warnings for each condition. Each condition had three warnings, one per email: IP address, Top-level domain mispositioned, and Link mismatch.
10.1 Claude feature-based explanations
@@
10.2 Llama feature-based explanations
@@
  • Link mismatch. The link https://www.facebook.com/hacked/disavow?u=100000125023309&nArdInDS2&lit_IT&ext1548538159 is an imitation of the original Facebook link. This site might be intended to take you to a different place. You might be disclosing private information.
10.3 Claude counterfactual explanations
@@
10.4 Llama counterfactual explanations
@@
  • Link mismatch. The link text “protect your account” hides the actual link https://www.facebook.com/hacked/disavow?u=100000125023309&nArdInDS2&lit_IT&ext1548538159. The email would have been considered safe if the link text matched the actual link, which helps ensure the link is trustworthy and not trying to trick you. A safe link might have looked like: https://www.facebook.com/account-security.
10.5 Baseline - Manually-generated feature-based explanations
@@
  • Top-level domain mispositioned. In the URL present in the email (https://amazonservices.com.cz/account.php), the top-level domain (e.g., “.com“) is in an abnormal position. This could indicate that the URL leads to a fake website. Such websites might steal your personal information.
  • Link mismatch. This email reports a link that is different from the actual one https://www.facebook.com/hacked/disavow?u=100000125023309&nArdInDS2&lit_IT&ext1548538159. This site might be intended to take you to a different place. You might be disclosing private information.
11 C. Logistic Regression Detailed Results
Table 4: Logistic regression results on CTR (click-through rate). OR>>1 increases click likelihood; OR<<1 decreases it.
Variable Cond. Emails OR $p$ Size
@@
Don’t continue CF agg ALL 0.21 <<.001 Large
CF agg TP 0.18 <<.001 Large
NoLLM ALL 0.14 <<.001 Large
NoLLM TP 0.10 <<.001 Large
@@
Avg hours/day NoLLM ALL 1.24 .024 Small
@@
Age × Hours FB agg ALL 1.007 .025 Negl.
CF agg ALL 1.007 .025 Negl.
Llama agg ALL 1.007 .025 Negl.
Claude agg ALL 1.007 .025 Negl.
NoLLM TP 1.02 .027 Negl.
Claude CF FP 0.93 .014 Negl.
Expertise × Hours Claude FB ALL 1.07 .042 Negl.
Llama CF ALL 1.11 .039 Negl.
Male × NFC CF agg FP 1.44 .001 Moderate
@@
Llama agg ALL 1.04 .032 Negl.
Expertise × NFC Llama agg ALL 1.03 .032 Negl.
Age × Expertise CF agg ALL 0.98 .048 Negl.
CF agg FP 0.97 .045 Negl.
Male × Age Llama agg TP 0.88 .02 Negl.
  • Effect sizes: Small (OR 1.2–1.5/0.67–0.83); Moderate (1.51–2/0.5–0.66); Large (>>2/<<0.5). ORs between 0.84–1.19 are Negligible (Negl.).